# User-Creator Feature Dynamics in Recommender Systems with Dual Influence

### v1

@@ 4 Discussions on Real-World Designs
The result in the previous section raises a natural question of whether polarization can be mitigated through careful design of recommender systems. We discuss 4 types of designs: top-$k$ recommendation, truncation, diversity-boosting, and uniform traffic.
(1) Top-$k$ Recommendation
A prevalent practice in modern two-stage recommendation algorithms on large-scale platforms, such as YouTube (Covington et al., 2016), is to first filter out items that are unlikely to be relevant to a user, then make recommendations from the remaining items. This practice reduces the computation cost and improves the relevancy of recommendations. Interestingly, we show that such a practice can also improve the long-term diversity of a recommender system with dual influence.
In particular, we consider the top-$k$ recommendation policy where each user is recommended only the $k$ most relevant creators, so $p_{ij}^{t}=0$ if $i$ is not one of $k$ creators $i^{\prime}$ that maximize $\langle\bm{v}_{i^{\prime}}^{t},\bm{u}_{j}^{t}\rangle$. We show that top-$k$ recommendation allows the dynamics to potentially converge to states with multiple clusters, instead of bi-polarization.
@@
With top-$k$ recommendation, there exist absorbing states that form $\lfloor n/k\rfloor$ clusters (i.e., once the system forms $\lfloor n/k\rfloor$ clusters, it forms $\lfloor n/k\rfloor$ clusters forever).
In contrast, Theorem 3.3 said that a recommender system where every creator can be recommended to every user ($p_{ij}^{t}>0$) is doomed to polarize. But with top-$k$ recommendation where some $p_{ij}^{t}=0$, polarization can be avoided. Experiments in Appendix B.3 also show that top-$k$ recommendation can reduce polarization and improve diversity.
(2) Truncation
Besides top-$k$ recommendation, truncation is another way to filter out irrelevant creators: set a threshold $\tau\in[-1,1]$ such that any user-creator pair with inner product $\langle\bm{u}_{i},\bm{v}_{j}\rangle<\tau$ is not recommended. Increasing $\tau$ is to similar to decreasing $k$ in top-$k$ recommendation, which improves recommendation relevance. Our experiments (in Appendix B) show that truncation at a large threshold helps to reduce polarization and improve diversity in a recommender system with dual influence.
@@
Diversity boosting tries to explore users’ interests and improve users’ experience by diversifying recommendation. For example, when making recommendations, the model optimizes the objective $h_{rel}(\langle\bm{u}_{i},\bm{v}_{j}\rangle)+\rho\cdot h_{div}(list_{i},\bm{v}_{j}),$ where $h_{rel},h_{div}$ rewards the recommendation relevance and diversity respectively and $list_{i}$ records the recent list of recommended items to user $i$. $h_{div}$ can take a simple form of $\sum_{j^{\prime}\in list_{i}}1-\langle\bm{v}_{j^{\prime}},\bm{v}_{j}\rangle$, and $\rho>0$ controls the strength of diversity-boosting. Despite being successful with fixed items and user preferences, this design cannot prevent bi-polarization in our dual-influence dynamics, since the conditions in Theorem 3.3 are still satisfied. Experiments in App. B.2 support our claim.
@@
Another method proposed in previous works (Judith Möller & van Es, 2018; Gao et al., 2022; Borgs et al., 2023) to improve the diversity of recommendation is to add a small fraction of uniform traffic to the personalized recommendation. As we argued in Sectin 3, this method could cause polarization and diversity loss in a recommender system with dual influence. Such an observation is striking as it demonstrates that optimizing for recommendation diversity in a static setting can ultimately lead to a huge loss of the system diversity in the long run.
5 Observations from Experiments
We conducted experiments on synthetic data and real-world data (MovieLens 20M (Harper & Konstan, 2015)). We simulated the user-creator feature dynamics for hundreds of time steps and observed the changes of four measures for the diversity of the system (Creator Diversity, Recommendation Diversity, Recommendation Relevance, and Tendency to Polarization) under different choices of parameters and real-world designs. Here are our three main observations:
  • Smaller sensitivity $\beta$ in the softmax function, larger creator update rates $\eta_{c}$, and larger user update rate $\eta_{u}$ cause faster polarization and worse diversity for the system. Notably, uniform recommendation ($\beta=0$), with the most diverse recommendation for the users in the short term, leads to severe polarization in the long run.
  • Top-$k$ recommendation with a small $k$ (Fig. 1) and truncation with a large threshold $\tau$ can alleviate polarization and improve the diversity of the system in the long run.
  • Diversity boosting methods, despite having a diversity-aware objective, fasten polarization and worsen the diversity of the system in the long run.
In summary, we found that, due to the dual influence of recommender systems, naïve diversity-promoting methods (like decreasing $\beta$ and using diversity-aware objective) cannot improve and even hurt the long-term diversity of the system; this is consistent with our Theorem 3.3 because those methods satisfy $p_{ij}^{t}\geq p_{0}>0$. Methods targeting efficiency and relevancy of recommendation (like top-$k$ recommendation), not satisfying $p_{ij}^{t}\geq p_{0}>0$, can instead improve the long-term diversity of the system.
Figure 1: Changes of diversity measures over time under top-$k$ recommendation different $k$, on MovieLens 20M dataset
6 Conclusion
We define a dynamics model to capture the dual influence of recommender systems on user preferences and content creation. Despite being a simplification, we believe that our model captures the essence of a real-world recommender system, and our effort is an important initial endeavor to study diversity in recommender systems with dual influence. We theoretically and empirically demonstrate that, due to dual influence, myopically optimizing recommendation diversity could cause polarization and long-term diversity loss. And somewhat unexpectedly, designs not targeting diversity (e.g., top-$k$ recommendation) can alleviate polarization. The insights from our work are valuable to building healthy and sustainable recommender systems, and we believe our results can inspire more sophisticated solutions for improving the diversity of recommender systems to be developed.
Appendix A Additional Discussion on Related Works
Diversity in recommendations
Diversity, filter bubbles, and polarization in recommendations have been important research topics in recent years, and they are closely related concepts with different focuses. On the one hand, filter bubbles are frequently defined as decreasing recommendation diversity over time (Aridor et al., 2020), which describes both the process and the outcome of insufficiently diverse recommendations. On the other hand, polarization describes the negative outcome of insufficient mutual understanding between people (Santos et al., 2021). In content platforms, an example of polarization is people creating content with strong agreement or disagreement with other content under the same topic, e.g., political opinions. To combat these negative outcomes, previous works propose diversity-boosting approaches including re-ranking (Carbonell & Goldstein, 1998; Ziegler et al., 2005) and diversity-aware objective optimization (Su et al., 2013; Zhang & Hurley, 2008; Hurley, 2013; Wilhelm et al., 2018; Cheng et al., 2017; Zhang et al., 2023). Despite having positive effects in situations where user preferences and creation styles are fixed, these approaches overlooked the dynamic nature of recommender systems and our work shows that certain approaches will make long-term outcomes worse under the dual influence.
Opinion dynamics
Opinion dynamics study the effect of people exchanging opinions with others on social networks (Sarlette et al., 2007; Golub & Jackson, 2010; Li & Spong, 2014; Altafini & Lini, 2015). Our model of a recommender system with dual influence on users and creators resembles a bipartite social network, and our conclusion that the system converges to polarization is conceptually similar to people reaching consensus on social networks (Acemoğlu et al., 2013; Caponigro et al., 2015; Markdahl et al., 2018; Zhang et al., 2022). However, the technique we use to prove our conclusion (absorbing Markov chain) significantly differs from the main technique (stability of ODE) in the mentioned works.
Performative prediction
Works in performative prediction assume that predictive systems (e.g., recommender systems) impact the individuals interacting with those systems (e.g., users and creators) (Perdomo et al., 2020; Hardt et al., 2022). These impacts can be direct, such as individuals ostensibly modifying their features in order to obtain more desirable outcomes (Levanon & Rosenfeld, 2022). Prior works on the performative effects of recommender systems (Eilat & Rosenfeld, 2023; Dean & Morgenstern, 2022) only consider one-sided impact, either on users or on creators. Differing from them, our work focuses on two-sided impacts, i.e., on both users and creators.
Appendix B Experiments
@@
We present experimental results on the behavior of the user-creator feature dynamics on synthetic data and real-world (MovieLens 20M) data and the effects of top-$k$ recommendation and truncation on the dynamics.
B.1 Synthetic Data Experiments
@@
The dynamics is initialized by randomly generating user and creator features on the unit sphere in ${\mathbb{R}}^{d}$. We pick $d=10$, number of creators $n=50$, number of users $m=100$. We use the softmax recommendation probability function (2). We simulate the dynamics for $T=1000$ steps, repeated $100$ times each with a new initialization. We choose the sign impact function $g(\bm{u}_{j},\bm{v}_{i})=\mathrm{sign}(\langle\bm{u}_{j},\bm{v}_{i}\rangle)$ for creator updates. For user updates, we choose inner product $f(\bm{v}_{i},\bm{u}_{j})=\langle\bm{v}_{i},\bm{u}_{j}\rangle$. The inner product function is studied in previous works on users’ preference dynamics (but not creators’) (Dean & Morgenstern, 2022). Note that the inner product does not satisfy the condition $|f(\bm{v}_{i},\bm{u}_{j})|\geq L_{f}$ needed in Theorem 3.3. However, we still observe convergence to polarization in nearly all experiments. Thus, even when this condition does not hold, users and creators still tend towards polarization in practice.
@@
  • Creator Diversity (CD): diversity of the creator features, measured by their average pairwise distance (Ziegler et al., 2005; Nguyen et al., 2014):
    $\mathrm{CD}(\bm{V})=\frac{1}{n(n-1)}\sum_{i=1}^{n}\sum_{j\neq i}\|\bm{v}_{i}-\bm{v}_{j}\|.$ (5)
  • Recommendation Diversity (RD): diversity of the contents recommended to a user, measured by the weighted variance of the contents:
    $\mathrm{RD}(\bm{U},\bm{V};\beta)=\frac{1}{m}\sum_{j=1}^{m}\sum_{i=1}^{n}p_{ij}\|\bm{v}_{i}-\overline{\bm{v}}_{j}\|^{2},$ (6)
    where $\overline{\bm{v}}_{j}=\sum_{i=1}^{n}p_{ij}\bm{v}_{i}$ and $p_{ij}=\frac{\exp(\beta\langle\bm{u}_{j},\bm{v}_{i}\rangle)}{\sum_{i=1}^{n}\exp(\beta\langle\bm{u}_{j},\bm{v}_{i}\rangle)}$.
  • Recommendation Relevance (RR): relevance of the contents recommended to a user, measured by the weighted average of inner products:
    $\mathrm{RR}(\bm{U},\bm{V};\beta)=\frac{1}{m}\sum_{j=1}^{m}\sum_{i=1}^{n}p_{ij}\langle\bm{u}_{j},\bm{v}_{i}\rangle.$ (7)
  • Tendency to Polarization (TP): This is a novel measure we propose to quantify how close the system is to consensus or bi-polarization, measured by the average absolute inner products between the creators:
    $\mathrm{TP}(\bm{V})=\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{k=1}^{n}|\langle\bm{v}_{i},\bm{v}_{k}\rangle|.$ (8)
    $\mathrm{TP}(\bm{V})$ being closer to 1 means that the system is more polarized, because the term $|\langle\bm{v}_{i},\bm{v}_{k}\rangle|$ is $1$ iff the two vectors $\bm{v}_{i},\bm{v}_{k}$ are equal or opposite to each other.
@@
Figure 2: Snapshots of the dynamics simulated with the same initialization but different recommendation sensitivity $\beta$.
A larger $\beta$ resulted in more clusters at time step $t=200$.
Figure 3: Changes of measures over time under different sensitivity parameter $\beta$, on synthetic data
@@
An explanation for our observation is the following: when $\beta$ is smaller, each user receives more uniform recommendations across all creators, so for different creators the sets of users recommended to those creators have larger intersections. Since the creator updates are based on the sets of recommended users, different creators will be moving towards more similar directions. This leads to faster polarization. One can also predict this observation from Theorem 3.3: when $\beta$ is large, the minimum recommendation probability $p_{0}$ of the softmax function tends to 00, so it might take a long time for the system to converge to polarization, while with a small $\beta$ the system polarizes quickly.
@@
We also consider the scenario where some dimensions of the user feature vectors are fixed features and thus not updated from round to round (e.g., age, gender), which is a realistic scenario. Formally, we fix the first $k\leq d$ dimensions. The remaining $d-k$ dimensions $\bm{u}_{j}^{t}[k+1:d]=(u_{j}^{t}[k+1],\ldots,u_{j}^{t}[d])$ are updated according to the following rule: $\bm{u}_{j}^{t+1}[k+1:d]=\|\bm{u}_{j}^{t}[k+1:d]\|\cdot\mathcal{P}\big{(}\bm{u}_{j}^{t}[k+1:d]+\eta_{u}f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})\bm{v}_{i}^{t}[k+1:d]\big{)}$. The multiplication by $\|\bm{u}_{j}^{t}[k+1:d]\|$ ensures unit norm $\|\bm{u}_{j}^{t+1}\|=1$. The effect of the number of fixed dimensions $k$ on the dynamics is shown in Figure 6. The main observation is: as the number of fixed dimensions increases, the diversity of the system improves and the degree of polarization is reduced. This is similar to the effect of decreasing user update rate $\eta_{u}$. The observation that fixed dimensions of user features help to improve diversity might be a reason why the recommender systems in practice are not as polarized as our theoretical prediction.
B.2 Real-World Data Experiments
Algorithm 1 Real-world Recommendation with Dual Influence
  repeat
     Let temporary parameter $\bm{w}^{(0)}\leftarrow\bm{\theta}^{(t)}$
     Compute loss $\mathcal{L}(\bm{\theta}^{(t)},Y^{(t)})$
     for $s=1$ to $m-1$ do
     end for
     $\bm{\theta}^{(t+1)}\leftarrow\bm{w}^{(m)}$
     Deliver recommendations based on $\hat{U}^{(t+1)},\hat{V}^{(t+1)}$
     Update $U^{(t+1)},V^{(t+1)}$, and $Y^{(t+1)}$
     $t\leftarrow t+1$
  until $\|\bm{\theta}^{(t)}-\bm{\theta}^{(t-1)}\|_{2}\leq\delta$
@@ B.2 Real-World Data Experiments
In this part, we conduct experiments on the MovieLens 20M dataset (Harper & Konstan, 2015). We use a real-world two-tower recommendation model with 16-dimensional tower tops as the user and creator embeddings (Fig. 7). The model was initialized on the original data, which we treat as the initial user and creator embeddings. Then we follow Algorithm 1 to simulate the dynamics.
Figure 9 shows the effect of using diversity-aware objective for diversity boosting (see Section 4). We see that myopically promoting the short-term recommendation diversity (using a larger $\rho$) results in a higher tendency to polarization and a lower creation diversity in the long run.
B.3 Top-$k$ Recommendation
We experimented with top-$k$ recommendation on the synthetic data and the MovieLens 20M dataset. Our main observation is: a small $k$ improves the diversity of the recommender system and reduces polarization. This is consistent with our theoretical prediction (Proposition 4.1). However, there is a tradeoff between the diversity of recommendations to users (RD) and the diversity of creations in the system (CD and TP). A top-$k$ recommendation policy with small $k$ is “not diverse” for users because it exposes a user only to a small set of contents. However, such a policy can lead to a more diverse outcome in the whole system. This tradeoff is worth further studying.
Results on the synthetic data set are in Table 1.
Table 1: Diversity improvement by top-$k$ recommendation on synthetic data

### v2

@@ 4 Discussions on Real-World Designs
Next, we discuss how 4 types of real-world recommender system designs affect the user-creator feature dynamics: top-$k$ truncation, threshold truncation, diversity-boosting, and uniform traffic.
(1) Top-$k$ Truncation
A prevalent practice in modern two-stage recommendation algorithms on large-scale platforms, such as YouTube [13], is to first filter out items that are unlikely to be relevant to a user, then make recommendations from the remaining items. In particular, we consider the top-$k$ truncation policy: for every user $j$, find the $k$ most relevant creators, namely, the $k$ creators whose inner products with the user $\langle\bm{v}_{i}^{t},\bm{u}_{j}^{t}\rangle$ are largest (equivalently, the $k$ creators whose probabilities $p_{ij}^{t}$ of being recommended to user $j$ are highest), then recommend one of those $k$ creators to user $j$ with probability proportional to $p_{ij}^{t}$. The other creators will not be recommended. This practice significantly reduces the computation cost and improves the relevancy of recommendations. Interestingly, we show that such a practice also has the potential to improve the long-term diversity of a recommender system with dual influence.
Definition 4.1 (clusters).
@@
With top-$k$ truncation, there exist states $(\bm{U}^{t},\bm{V}^{t})$ that form $\lfloor n/k\rfloor$ clusters and are absorbing (i.e., once the system forms $\lfloor n/k\rfloor$ clusters, it forms $\lfloor n/k\rfloor$ clusters forever).
This result is in contrast with Theorem 3.3 which shows that a recommender system where every creator can be recommended to every user ($p_{ij}^{t}>0$) is doomed to polarize. With top-$k$ truncation where some $p_{ij}^{t}=0$, polarization can be avoided. Experiments in Section 5.3 support our prediction that top-$k$ truncation can reduce polarization and improve diversity.
(2) Threshold Truncation
Besides top-$k$ truncation, threshold truncation is another way to filter out irrelevant creators: set a threshold $\tau\in[-1,1]$ such that any user-creator pair with inner product $\langle\bm{u}_{i},\bm{v}_{j}\rangle<\tau$ is not recommended. A natural choice is $\tau=0$, meaning that users will not receive recommendations predicted to be “disliked” by them. Increasing $\tau$ is similar to increasing the $\beta$ in the softmax function, which improves recommendation relevance.
Proposition 4.3.
In $d$-dimensional feature space, if user-creator pairs with $\langle\bm{u}_{i},\bm{v}_{j}\rangle<0$ are not recommended, then there exist stable states with $d+1$ clusters.
Although truncation at $\tau=0$ allows stable states with $d+1$ clusters to exist, the dynamics does not necessarily converge to such states; it can still end up with stable states with fewer clusters. In fact, experiments (in Appendix B.2) show that truncation at $\tau=0$ is not good for diversity and causes severe polarization, while truncation at a large threshold like $\tau=0.707$ is better at reducing polarization.
@@
Diversity boosting aims to explore users’ interests and improve users’ experience by diversifying recommendation. For example, when making recommendations, the model optimizes the objective:
$h_{rel}(\langle\bm{u}_{i},\bm{v}_{j}\rangle)+\rho h_{div}(list_{i},\bm{v}_{j}),$ (5)
where $h_{rel},h_{div}$ rewards the recommendation relevance and diversity respectively and $list_{i}$ records the recent list of recommended items to user $i$. $h_{div}$ can take a simple form of $\sum_{j^{\prime}\in list_{i}}1-\langle\bm{v}_{j^{\prime}},\bm{v}_{j}\rangle$, and $\rho>0$ controls the strength of diversity-boosting. Despite being successful when users’ preferences and items are fixed, this design alone cannot prevent bi-polarization in our dual-influence dynamics, since the conditions in Theorem 3.3 are still satisfied and the users’ and creators’ update rules remain the same. Experiments in Section 5.2 support our claim.
@@
Adding a small fraction of uniform traffic to the personalized recommendations is another method proposed in previous works to improve recommendation diversity or to explore user preferences [25, 16, 9, 8, 30]. This method gives a non-zero lower bound on the probability of every creator being recommended to every user. So, as a corollary of our Theorem 3.3, it causes a recommender system with dual influence to polarize. Such an observation is striking as it demonstrates that optimizing for recommendation diversity in a static setting can ultimately lead to a huge loss of the system diversity in the long run.
5 Experiments
@@
We present experimental results on the behavior of user-creator feature dynamics on synthetic data and real-world (MovieLens 20M) data and the effect of top-$k$ truncation and threshold truncation on the dynamics.
5.1 Synthetic Data Experiments
@@
The dynamics is initialized by randomly generating user and creator features on the unit sphere in ${\mathbb{R}}^{d}$. We pick $d=10$, number of creators $n=50$, number of users $m=100$. We use the softmax recommendation probability function (2). We simulate the dynamics for $T=1000$ steps, repeated $100$ times each with a new initialization. We choose the sign impact function $g(\bm{u}_{j},\bm{v}_{i})=\mathrm{sign}(\langle\bm{u}_{j},\bm{v}_{i}\rangle)$ for creator updates. For user updates, we choose inner product $f(\bm{v}_{i},\bm{u}_{j})=\langle\bm{v}_{i},\bm{u}_{j}\rangle$. The inner product function is studied in previous works on users’ preference dynamics (but not creators’) [14]. Note that the inner product does not satisfy the condition $|f(\bm{v}_{i},\bm{u}_{j})|\geq L_{f}$ needed in Theorem 3.3. However, we still observe convergence to polarization in nearly all experiments. Thus, even when this condition does not hold, users and creators still tend towards polarization in practice.
@@
  • Creator Diversity (CD): diversity of the creator features, measured by their average pairwise distance [47, 33]: $\mathrm{CD}(\bm{V})=\frac{1}{n(n-1)}\sum_{i=1}^{n}\sum_{j\neq i}\|\bm{v}_{i}-\bm{v}_{j}\|$.
  • Recommendation Diversity (RD): diversity of the contents recommended to a user, measured by the weighted variance of the contents: $\mathrm{RD}(\bm{U},\bm{V};\beta)=\frac{1}{m}\sum_{j=1}^{m}\sum_{i=1}^{n}p_{ij}\|\bm{v}_{i}-\overline{\bm{v}}_{j}\|^{2}$, where $\overline{\bm{v}}_{j}=\sum_{i=1}^{n}p_{ij}\bm{v}_{i}$ and $p_{ij}=\frac{\exp(\beta\langle\bm{u}_{j},\bm{v}_{i}\rangle)}{\sum_{i=1}^{n}\exp(\beta\langle\bm{u}_{j},\bm{v}_{i}\rangle)}$.
  • Recommendation Relevance (RR): relevance of the contents recommended to a user, measured by the weighted average of inner products: $\mathrm{RR}(\bm{U},\bm{V};\beta)=\frac{1}{m}\sum_{j=1}^{m}\sum_{i=1}^{n}p_{ij}\langle\bm{u}_{j},\bm{v}_{i}\rangle$.
  • Tendency to Polarization (TP): This is a novel measure we propose to quantify how close the system is to consensus or bi-polarization, measured by the average absolute inner products between the creators: $\mathrm{TP}(\bm{V})=\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{k=1}^{n}|\langle\bm{v}_{i},\bm{v}_{k}\rangle|$. $\mathrm{TP}(\bm{V})$ being closer to 1 means that the system is more polarized, because the term $|\langle\bm{v}_{i},\bm{v}_{k}\rangle|$ is $1$ iff the two vectors $\bm{v}_{i},\bm{v}_{k}$ are equal or opposite to each other.
@@
Figure 1: Snapshots of the dynamics simulated with the same initialization but different recommendation sensitivity $\beta$. A larger $\beta$ resulted in more clusters at time step $t=200$.
Figure 2: Changes of measures over time under different sensitivity parameter $\beta$, on synthetic data. $\beta=0$ means uniform (non-personalized) recommendation. $\beta=\infty$ means hard-max recommendation: only recommend the single most relevant creator to a user. Larger $\beta$ reduces the tendency to polarization.
@@
An explanation for our observation is the following: When $\beta$ is smaller, each user receives more uniform recommendations across all creators. So, for different creators, the sets of users recommended to those creators have larger intersections. Since the creator updates are based on the sets of recommended users, different creators will be moving towards more similar directions. This leads to faster polarization. One can also predict this observation from Theorem 3.3: when $\beta$ is large, the minimum recommendation probability $p_{0}$ of the softmax function tends to 00, so it might take a long time for the system to converge to polarization, while with a small $\beta$ the system polarizes quickly.
@@
We also consider the scenario where some dimensions of the user feature vectors are fixed features and thus not updated from round to round (e.g., age, gender), which is a realistic scenario. Detailed results are in Appendix B.1. The main observation is: as the number of fixed dimensions increases, the diversity of the system improves and the degree of polarization is reduced. This is similar to the effect of decreasing user update rate $\eta_{u}$. The observation that fixed dimensions of user features help to improve diversity might be a reason why the recommender systems in practice are not as polarized as our theoretical prediction.
5.2 Real-World Data Experiments
In this part, we conduct experiments on the MovieLens 20M dataset [19]. We use a real-world two-tower recommendation model with 16-dimensional tower tops as the user and creator embeddings (Figure 9). The model is initialized by fitting a two-tower model [21] on the existing MovieLens rating data and using the tower tops as the initial user and creator embeddings. Then we follow Algorithm 1 to simulate the dynamics.
Figure 6 shows the effect of using diversity-aware objective (Eq. 5) for diversity boosting. We see that myopically promoting the short-term recommendation diversity (using a larger $\rho$) results a higher creation diversity but also a higher tendency to polarization. A possible explanation for this phenomenon is, similar to the case with $\beta$, the system polarizes into two balanced clusters which actually have a large average pairwise distance. In this case, Tendency to Polarization is a better measure for diversity loss than Creator Diversity (average pairwise distance).
@@ 5.2 Real-World Data Experiments
5.3 Top-$k$ Truncation and Threshold Truncation
We experimented with top-$k$ truncation on the synthetic data (Table 1) and the MovieLens dataset (Appendix C). Our main observation is: a small $k$ improves the diversity of the recommender system and reduces polarization. This is consistent with our theoretical prediction (Proposition 4.2). However, there is a tradeoff between the diversity of recommendations to users (RD) and the diversity of creations in the system (CD and TP). A top-$k$ truncation policy with small $k$ is “not diverse” for users because it exposes a user only to a small set of contents. However, such a policy can lead to a more diverse outcome in the whole system. This tradeoff is worth further studying.
Table 1: Diversity improvement by top-$k$ truncation on synthetic data